### v2

@@ Abstract
Dairy farming consumes a significant amount of energy, making it an energy-intensive sector within agriculture. Integrating renewable energy generation into dairy farming could help address this challenge. Effective battery management is important for integrating renewable energy generation. Managing battery charging and discharging poses significant challenges because of fluctuations in electrical consumption, the intermittent nature of renewable energy generation, and fluctuations in energy prices. Artificial Intelligence (AI) has the potential to significantly improve the use of renewable energy in dairy farming, however, there is limited research conducted in this particular domain. This research considers Ireland as a case study as it works towards attaining its 2030 energy strategy centered on the utilization of renewable sources. This study proposes a Q-learning-based algorithm for scheduling battery charging and discharging in a dairy farm setting. This research also explores the effect of the proposed algorithm by adding wind generation data and considering additional case studies. The proposed algorithm reduces the cost of imported electricity from the grid by 13.41%, peak demand by 2%, and 24.49% when utilizing wind generation. These results underline how reinforcement learning is highly effective in managing batteries in the dairy farming sector.
@@ 1 Introduction
The growth in the global population has led to an increased demand for food products. Milk maintains an important place in the dietary patterns of individuals across the globe because of its essential nutritious components. According to the Food and Agriculture Organisation (FAO), global milk production rose from 735 million metric tonnes in 2000 to 855 million metric tonnes in 2019 b2 . The rising demand for dairy products has led to an increase in the number of dairy farmsbib_oecd2020dairy . These farms heavily rely on electricity for multiple activities like milk cooling, water heating, pumping, and lighting b3 . Meeting these energy needs requires substantial imports of electricity from external grids. However, the rising cost of electricity necessitates considering alternative energy sources like solar photovoltaic and wind turbines. Embracing renewable energy sources can help satisfy the energy requirements of farms and decrease dependence on the external grid for importing electricity b4 . By 2030 Irish government aims to transition to a low-carbon economy within the EU, emphasizing renewable energy, secure electricity supply, and enhanced energy efficiency energy . This research supports these goals by efficiently managing energy storage, increasing renewable energy use, and reducing carbon emissions by minimizing grid reliance as the grid generates most of the energy by burning fossil fuels grid_gen .
@@
Power generated from renewable energy sources exhibits temporal variability. Employing a battery management system for storing electrical energy is crucial for future use. Different battery management systems are applied to different applicationshannan2021battery . The use of batteries has the potential to influence the economic aspects of electricity consumption within dairy farming. However, optimizing battery performance necessitates the implementation of different strategies. Various conventional methods have been employed to improve battery efficiency. These approaches include such as Maximising Self-Consumption (MSC) and Time of Use (TOU) b5 .
@@
In recent years, there has been remarkable progress in Artificial Intelligence (AI), largely driven by the data revolution. This advancement has shown immense potential across various fields, yielding promising resultsai_application . One significant area within AI is Reinforcement Learning (RL), where agents can operate in stochastic environments without explicit knowledge of the environment or predefined decision-making processes. Instead, with established objective functions. Two main algorithms that are particularly prominent in the field of RL are actor-critic learning and Q-learning acl ; b12 . RL agents can effectively learn a policy in diverse domains through these algorithms. This research aims to use RL agents to learn battery management policies efficiently. The primary goal is to optimize the charging/discharging of a battery to maximize renewable utilization and reduce the cost of electricity imported from the grid.
Previous studies have shown the effectiveness of RL in managing batteries across various settings. Ebell et al. utilized RL methods to control a battery system in household settings equipped with photovoltaic (PV) panels. Their findings demonstrated a 7.8% reduction in grid energy usage b8 . However, implementing RL algorithms for battery management within dairy farms improves the effectiveness of battery management. Integrating RL into the agricultural sector, particularly in the dairy farming industry, remains a topic that has received limited attention in current literature.
The main contributions of this research are specified below:
@@
    To apply Q-Learning to dairy farm battery management.
@@
    To investigate the impact of wind generation on battery management.
@@
    To evaluate how different state space information affects Q-Learning.
@@
    To evaluate the performance of Q-Learning using Finland and Ireland case studies.
@@ 2 Background and Related Work
2.1 Conventional Battery Control Methods
Many studies have focused on identifying effective operational strategies for PV battery systems, for different objectives b13 . Specifically, the Maximising Self-Consumption (MSC) and Time of Use (TOU) methods for battery charging b14 .
The MSC is a method used for managing battery charging and discharging by maximizing the utilization of solar power generation. It charges the maximum amount of solar energy available b13 . Braun et al. highlighted that optimal battery usage significantly increases the local consumption of solar energyb15 . Further, they investigated the optimal sizing of photovoltaic systems b16 and the capacity requirements for energy storage b17 , aiming to maximize the use of locally generated solar power and reduce reliance on external power grids. In their comprehensive review, Luthander et al. analyzed previous studies on solar power self-consumption in buildings, concluding that proper battery sizing can improve self-consumption rates by 13-24% b18 . Sharma et al. conducted a study on the optimization of battery size for zero-net energy homes equipped with rooftop solar panels in South Australia, employing the MSC operational strategy b19 . Their findings suggest that installing suitable batteries can enhance the self-consumption of solar energy by 20-50% b20 .
TOU strategy uses electricity prices for charging and discharging the battery; it charges the battery when the prices are low and discharges at peak times. Feed-in Tariffs (FiT) and TOU pricing strategies, implemented in several countries, aim to enhance the adoption of Photovoltaic Battery (PVB) systems and encourage consumer involvement in energy management, which has been a significant area of research. b21 . Other studies have focused on the TOU tariff method for efficient battery management. For instance, Christoph et al. utilized optimization techniques to refine the TOU rate structure b22 , while Li et al. developed TOU tariffs using the Gaussian Mixture Model b23 . This approach has enabled prosumers to get economic advantages by taking advantage of FiT and adapting to varying electricity prices during peak and off-peak times, which is a main benefit of the TOU strategy b24 . Research by Gitizadeh et al. and Hassan et al. explores optimizing battery capacity, by utilizing TOU b25 b26 . Additionally, Ratnam et al. found that many PVB system users were able to achieve significant annual cost reductions through FiT programs b27 .
@@ 2.2 Reinforcement Learning
[…]
@@ 2.3 Reinforcement Learning for Energy Management
RL algorithms are widely used in various applications. Wei et al. implemented dual iterative Q-learning for managing batteries in smart residential settings b28 . This form of Q-learning is designed to enhance energy management in smart homes by optimizing the charging and discharging of batteries. Similarly, Kim et al. developed an RL-based algorithm for energy management in smart buildings b29 . Their approach uses RL to dynamically identify the most effective energy regulation strategy based on real-time data. Ruelens et al. also applied RL, but to the operation of an electric water heater b30 , using the algorithm to boost the heater’s energy efficiency by learning and adapting to real-time user demand and grid conditions. Research indicates that RL can significantly enhance both efficiency and cost-effectiveness in energy consumption within smart grids. Furthermore, Li et al. introduced a multi-grid RL method to optimize the energy efficiency and comfort of Heating, Ventilation, and Air Conditioning (HVAC) systems b31 . This method balances HVAC energy use with maintaining optimal room temperature and humidity. Their findings suggest that this approach effectively optimizes energy consumption while ensuring comfortable indoor environments.
@@ 2.4 Reinforcement Learning for Battery Management
Numerous studies have explored battery management using RL. However, none of these studies have explicitly focused on its implications within the context of dairy farming. Foruzan et al. introduced the use of RL for managing energy in microgrids b32 . They employed an RL system capable of adapting in real-time to changing energy needs and the generation of renewable energy, enhancing the energy efficiency of microgrids. RL is effective in improving energy consumption in a cost-efficient manner. In a similar application, Guan et al. developed an RL-based solution for controlling domestic energy storage to reduce electricity cost b33 . This RL method optimizes the charging and discharging of energy storage systems, helping to decrease peak power demands and shift energy usage to cheaper, off-peak times, thereby lowering electricity bills. Their simulations demonstrated that this strategy could effectively reduce the electricity cost associated with household energy storage systems. Liu et al. explored the use of Deep Reinforcement Learning (DRL) for optimizing energy management in households b34 . This study utilized a DRL system designed to enhance energy efficiency in smart homes by constantly learning the most effective energy management strategies. In simulated smart home environments, this DRL-based approach was found to be more efficient and cost-effective than traditional rule-based methods, indicating its potential to significantly improve energy management in intelligent residential settings. Cao et al. proposed the DRL method for battery charging and discharging, handling power price uncertainty, improving the accuracy of the degradation model, and non-linear charging and discharging efficiency cao2020deep . They demonstrated the algorithm’s efficacy and performance by testing it on historical wholesale electricity data from the United Kingdom. Yu et al. use Deep Deterministic Policy Gradient (DDPG) for the home energy management system to minimize electricity cost by scheduling HVAC systems and Effective Solutions for Storing (ESSs) yu2019deep . By leveraging the dynamic prices, the results demonstrated that the proposed algorithm saves energy cost by 8.1%-15.21%.
@@
Abedi et al. have created a real-time intelligent battery energy control system for residential buildings that incorporates solar panels, battery energy systems, and grid connectivity by using Q-learning. The results of their study demonstrate that the algorithm effectively decreases the monthly electricity cost by 7.99% to 3.63% for house 27 and 6.91% to 3.26% for house 387. Wei et al. proposed the DDPG a DRL algorithm for the fast charging of lithium-ion batteries (LIB) wei2021deep . They compare the proposed algorithm with the rule base by considering different constraints i.e. LIB temperature, charging rapidity, and degradation. Huang et al. introduce Proximal Policy Optimisation (PPO) as a DRL algorithm to optimize the capacity scheduling of solar battery systems huang2020deep . To enhance the safety of the battery, a safety control algorithm is implemented by utilizing a serial approach incorporated with a PPO algorithm. Their findings indicate that the proposed algorithm outperforms other DRL algorithms. Cheng et al. propose a periodic deterministic policy gradient (PDPG) to schedule the charging of multi-battery energy storage systems (MBESS) cheng2023reinforcement . Their research shows that compared to the DPG algorithm, the PDPG algorithm reduces power cost by 8.79%. Paudel et al. employ the MDP framework to efficiently manage the charging and discharging operations of battery storage systems, by considering the electricity price fluctuations and other relevant parameters paudel2023deep . The authors substantiate their method’s effectiveness by installing 150 fast charging stations and a battery storage system throughout the Pennsylvania-New Jersey-Maryland region. The studies mentioned above show the impact of RL on the application of battery management. This research utilizes the Q-learning algorithm for battery management in dairy farming.
@@ 3.1 System Design
The PVB system connected to the grid, as depicted in Figure 1, includes a set of components: solar panels, a battery storage unit, the power grid, and a dairy farm that utilizes electricity from both solar and grid sources. The energy storage system considered for this research is the Tesla Powerwall 2.0, which offers a substantial capacity of 13.5kWh and supports both charging and discharging 5kW b36 . The PV-generated electricity is used to meet the farm’s load, charge the battery, or sell it back to the grid, according to the operational requirements. The role of the charge/discharge controller is to charge and discharge the battery according to the renewable generation, electricity demand, and price of electricity. Meanwhile, the power grid is connected to the dairy farm and the battery. It supplies electricity when there is high demand and low renewable generation. The battery storage is used to satisfy the farm’s extra energy needs, a process commonly referred to as peak shaving. This involves using excess energy demands by utilizing stored power in the batteryhannanpeakshaving .
@@ 3.2 Data and Price Profile
For this study, two datasets were used: One dataset from Finland to train the algorithm and a second dataset collected from Ireland to evaluate the performance of the algorithm. The Finland dataset has information about the load demand from dairy farms, PV generation, wind generation, and electricity prices. The load data is collected from b10 and provides hourly electricity consumption over a year. Figure 1(a) demonstrates the monthly distribution of electricity demands for a dairy farm and PV generation and wind energy generated by the dairy farm throughout the year. The dataset consists of a dairy farm that has approximately 180 cows and has an estimated annual electricity usage of around 261 megawatt-hours (MWh). The PV and wind data was collected from the System Advisor Model (SAM) having a capacity of 20kW b35 . The Finland electricity price data was collected from a Helsinki electricity supply company b_fin_prc . This price data is dynamic and includes three different price levels b_tou . The lowest rate is during off-peak hours, the standard rate applies for most of the day, and a higher peak rate is charged during the busiest hours. Specifically, the pricing is segmented into three time periods. The off-peak hours, with the lowest rate, are from 11 p.m. to 7 a.m. The standard rate applies during two intervals: from 8 a.m. to 4 p.m. and from 7 p.m. to 10 p.m. The peak rate, which is the highest, is charged between 5 p.m. and 7 p.m. Figure 1(b) shows how these electricity prices fluctuate over the day.
@@
The Ireland dataset includes data on farm load, PV generation, and electricity pricing. However, it lacks wind generation data as we were unable to find wind generation data for Ireland. The load consumption data, detailing electricity from the dairy farm over a year, was collected from a study on Irish dairy farms b_hu . The PV generation data was collected from SAMb35 having a capacity of 20kW. The price data is collected from the Ireland electricity supply company Electric Ireland b_ir_prc . Figure 3 shows the Irish dairy farm energy consumption and photovoltaic (PV) energy generation and electricity price. This figure illustrates the variations in PV generation and electricity price, it also demonstrates the farm electricity demand patterns. The aim is to explore the relationship between energy consumption and PV generation, particularly in the Irish dairy farm context. Figure 2(a) specifically illustrates the monthly load demand and PV generation of the dairy farm over one year, while Figure 2(b) illustrates the price variations over the day.

### v3

@@ Abstract
Dairy farming consumes a significant amount of energy, making it an energy-intensive sector within agriculture. Integrating renewable energy generation into dairy farming could help address this challenge. However, fluctuations in renewable generation pose a challenge to this integration. Effective battery management techniques are needed to store and utilize the energy generated from renewable sources. The objective of this research is to leverage Reinforcement Learning to develop an effective approach for battery management systems in dairy farming. Our work contributes by implementing a Q-learning algorithm for dairy farm battery management, incorporating wind and solar energy, exploring the state space of the algorithm, and considering Ireland as a case study as it works towards attaining its 2030 energy strategy centered on the utilization of renewable sources. The findings show that the proposed algorithm reduces the cost of imported electricity from the grid by 13.41%, 24.49% when utilizing wind generation, and peak demand by 2%. These findings highlight the effectiveness of Reinforcement Learning for battery management in the dairy farming sector.
@@ 1 Introduction
The growth in the global population has led to an increased demand for food products. Milk maintains an important place in the dietary patterns of individuals across the globe because of its essential nutritious components. According to the Food and Agriculture Organisation (FAO), global milk production rose from 735 million metric tonnes in 2000 to 855 million metric tonnes in 2019 [1]. The rising demand for dairy products has led to an increase in the number of dairy farms[2]. These farms heavily rely on electricity for multiple activities like milk cooling, water heating, pumping, and lighting [3]. Meeting these energy needs requires substantial imports of electricity from external grids. However, the rising cost of electricity necessitates considering alternative energy sources like solar photovoltaic and wind turbines. Embracing renewable energy sources can help satisfy the energy requirements of farms and decrease dependence on the external grid for importing electricity [4]. By 2030 Irish government aims to transition to a low-carbon economy within the EU, emphasizing renewable energy, secure electricity supply, and enhanced energy efficiency [5]. This research supports these goals by efficiently managing energy storage, increasing renewable energy use, and reducing carbon emissions by minimizing grid reliance as the grid generates most of the energy by burning fossil fuels [6].
@@
Power generated from renewable energy sources exhibits temporal variability. Employing a battery management system for storing electrical energy is crucial for future use. Different battery management systems are applied to different applications[7]. The use of batteries has the potential to influence the economic aspects of electricity consumption within dairy farming. However, optimizing battery performance necessitates the implementation of different strategies. Various conventional methods have been employed to improve battery efficiency. These approaches include such as Maximising Self-Consumption (MSC) and Time of Use (TOU) [8].
@@
In recent years, there has been remarkable progress in Artificial Intelligence (AI), largely driven by the data revolution. This advancement has shown immense potential across various fields, yielding promising results[9]. One significant area within AI is Reinforcement Learning (RL), where agents can operate in stochastic environments without explicit knowledge of the environment or predefined decision-making processes. Instead, with established objective functions. Two main algorithms that are particularly prominent in the field of RL are actor-critic learning and Q-learning [10, 11]. RL agents can effectively learn a policy in diverse domains through these algorithms. This research aims to use RL agents to learn battery management policies efficiently. The primary goal is to optimize the charging/discharging of a battery to maximize renewable utilization and reduce the cost of electricity imported from the grid.
Existing research has highlighted the efficacy of RL in battery management across diverse contexts. However, the integration of RL into the agricultural sector, specifically in dairy farming for battery management, remains relatively unexplored in current literature. The main contributions of this research are specified below:
@@
    Using Q-learning we present an autonomous learning approach for efficient battery management in dairy farms, and we demonstrate the effectiveness of our algorithm in achieving improved energy efficiency over established baseline algorithms.
@@
    In contrast to existing approaches, this work also analyses the influence of incorporating wind energy data on the effectiveness of battery management.
@@
    We evaluate the impact of variations in state space information on the performance of our Q-learning approach. We explore the impact of additional parameters including load, PV, and wind to determine the most optimal solution.
@@
    We extend our experimental analysis to also evaluate the Q-Learning algorithm’s performance using data based on case studies in Finland and Ireland, focusing on dairy farm battery management.
@@ 2.1 Reinforcement Learning
The update rule for the Q-value of the current state-action pair $(s_{t},a_{t})$ involves a scalar factor $\alpha$, Also known as a learning rate which controls the rate at which the agents will explore the environment ranging from 0 to 1 and scales the difference between the observed reward Rt plus the discounted estimate of the maximum Q-value for the next state $s_{t+1}$ (discounted by a factor $\gamma)$ and the current estimate of the Q-value for the current state-action pair $Q(s_{t},a_{t})$.
2.2 Conventional Battery Control Methods
Many studies have focused on identifying effective operational strategies for PV battery systems, for different objectives [12]. Specifically, the Maximising Self-Consumption (MSC) and Time of Use (TOU) methods for battery charging [13].
The MSC is a method used for managing battery charging and discharging by maximizing the utilization of solar power generation. It charges the maximum amount of solar energy available [12]. Braun et al. highlighted that optimal battery usage significantly increases the local consumption of solar energy[14]. Further, they investigated the optimal sizing of photovoltaic systems [15] and the capacity requirements for energy storage [16], aiming to maximize the use of locally generated solar power and reduce reliance on external power grids. In their comprehensive review, Luthander et al. analyzed previous studies on solar power self-consumption in buildings, concluding that proper battery sizing can improve self-consumption rates by 13-24% [17]. Sharma et al. conducted a study on the optimization of battery size for zero-net energy homes equipped with rooftop solar panels in South Australia, employing the MSC operational strategy [18]. Their findings suggest that installing suitable batteries can enhance the self-consumption of solar energy by 20-50% [19].
TOU strategy uses electricity prices for charging and discharging the battery; it charges the battery when the prices are low and discharges at peak times. Feed-in Tariffs (FiT) and TOU pricing strategies, implemented in several countries, aim to enhance the adoption of Photovoltaic Battery (PVB) systems and encourage consumer involvement in energy management, which has been a significant area of research. [20]. Other studies have focused on the TOU tariff method for efficient battery management. For instance, Christoph et al. utilized optimization techniques to refine the TOU rate structure [21], while Li et al. developed TOU tariffs using the Gaussian Mixture Model [22]. This approach has enabled prosumers to get economic advantages by taking advantage of FiT and adapting to varying electricity prices during peak and off-peak times, which is a main benefit of the TOU strategy [23]. Research by Gitizadeh et al. and Hassan et al. explores optimizing battery capacity, by utilizing TOU [24][25]. Additionally, Ratnam et al. found that many PVB system users were able to achieve significant annual cost reductions through FiT programs [26].
@@ 2.3 Reinforcement Learning for Energy Management
RL algorithms are widely used in various applications. Wei et al. implemented dual iterative Q-learning for managing batteries in smart residential settings [27]. This form of Q-learning is designed to enhance energy management in smart homes by optimizing the charging and discharging of batteries. Similarly, Kim et al. developed an RL-based algorithm for energy management in smart buildings [28]. Their approach uses RL to dynamically identify the most effective energy regulation strategy based on real-time data. Ruelens et al. also applied RL, but to the operation of an electric water heater [29], using the algorithm to boost the heater’s energy efficiency by learning and adapting to real-time user demand and grid conditions. Research indicates that RL can significantly enhance both efficiency and cost-effectiveness in energy consumption within smart grids. Furthermore, Li et al. introduced a multi-grid RL method to optimize the energy efficiency and comfort of Heating, Ventilation, and Air Conditioning (HVAC) systems [30]. This method balances HVAC energy use with maintaining optimal room temperature and humidity. Their findings suggest that this approach effectively optimizes energy consumption while ensuring comfortable indoor environments.
@@ 2.4 Reinforcement Learning for Battery Management
Numerous studies have explored battery management using RL. Foruzan et al. introduced the use of RL for managing energy in microgrids [31]. They employed an RL system capable of adapting in real-time to changing energy needs and generating renewable energy, enhancing the energy efficiency of microgrids. RL is effective in improving energy consumption in a cost-efficient manner. In a similar application, Guan et al. developed an RL-based solution for controlling domestic energy storage to reduce electricity cost [32]. This RL method optimizes the charging and discharging of energy storage systems, helping decrease peak power demands and shift energy usage to cheaper, off-peak times, lowering electricity bills. Their simulations demonstrated that this strategy could effectively reduce the electricity cost associated with household energy storage systems. Liu et al. explored the use of Deep Reinforcement Learning (DRL) for optimizing energy management in households [33]. This study utilized a DRL system designed to enhance energy efficiency in smart homes by constantly learning the most effective energy management strategies. In simulated smart home environments, this DRL-based approach was more efficient and cost-effective than traditional rule-based methods, indicating its potential to significantly improve energy management in intelligent residential settings. Cao et al. proposed the DRL method for battery charging and discharging, handling power price uncertainty, improving the accuracy of the degradation model, and non-linear charging and discharging efficiency [34]. They demonstrated the algorithm’s efficacy and performance by testing it on historical wholesale electricity data from the United Kingdom. Yu et al. use Deep Deterministic Policy Gradient (DDPG) for the home energy management system to minimize electricity cost by scheduling HVAC systems and Effective Solutions for Storing (ESSs) [35]. By leveraging the dynamic prices, the results demonstrated that the proposed algorithm saves energy cost by 8.1%-15.21%.
@@
Abedi et al. have created a real-time intelligent battery energy control system for residential buildings that incorporates solar panels, battery energy systems, and grid connectivity by using Q-learning. The results of their study demonstrate that the algorithm effectively decreases the monthly electricity cost by 7.99% to 3.63% for house 27 and 6.91% to 3.26% for house 387. Wei et al. proposed the DDPG a DRL algorithm for the fast charging of lithium-ion batteries (LIB) [36]. They compare the proposed algorithm with the rule base by considering different constraints i.e. LIB temperature, charging rapidity, and degradation. Huang et al. introduce Proximal Policy Optimisation (PPO) as a DRL algorithm to optimize the capacity scheduling of solar battery systems [37]. To enhance the safety of the battery, a safety control algorithm is implemented by utilizing a serial approach incorporated with a PPO algorithm. Their findings indicate that the proposed algorithm outperforms other DRL algorithms. Cheng et al. propose a periodic deterministic policy gradient (PDPG) to schedule the charging of multi-battery energy storage systems (MBESS) [38]. Their research shows that compared to the DPG algorithm, the PDPG algorithm reduces power cost by 8.79%. Paudel et al. employ the MDP framework to efficiently manage battery storage systems’ charging and discharging operations by considering the electricity price fluctuations and other relevant parameters [39]. The authors substantiate their method’s effectiveness by installing 150 fast charging stations and a battery storage system throughout the Pennsylvania-New Jersey-Maryland region. The studies mentioned above show RL’s impact on battery management applications.
The conventional and RL studies underscore the importance of maximizing local energy utilization and optimizing battery usage. However, some limitations have been identified in these works that our research aims to address. Firstly, they did not address performance variations under diverse weather conditions and geographical locations, besides the impact of fluctuating energy prices and renewable generation. Secondly, these studies focus solely on one renewable source and do not consider the effects of integrating other energy sources. Lastly, all the conventional and RL methods have been applied in smart homes and buildings, but their adaptation to dairy farm battery management remains largely unexplored. Dairy farms typically consume more energy than households or offices due to operational needs and reliance on high-energy equipment like milking machines and milk cooling systems, which account for 20-30% of the farm’s electricity. Furthermore, research has shown that electricity consumption per dairy cow ranges from 4 to 7.3 kWh/week[3]. In contrast, households and offices use energy mainly for heating, cooling, lighting, and appliances. The unique requirements of dairy farming operations lead to higher load consumption and diverse consumption patterns. This research addresses these gaps by demonstrating how a Q-learning algorithm optimizes battery management in dairy farming settings. It also mitigates drawbacks by testing the proposed methodology across various locations and weather conditions, integrating multiple renewable sources, and considering electricity price fluctuations.
@@ 3.1 System Design
The PVB system connected to the grid, as depicted in Figure 1, includes a set of components: solar panels, a battery storage unit, the power grid, and a dairy farm that utilizes electricity from both solar and grid sources. The energy storage system considered for this research is the Tesla Powerwall 2.0, which offers a substantial capacity of 13.5kWh and supports both charging and discharging 5kW [40]. The PV-generated electricity is used to meet the farm’s load, charge the battery, or sell it back to the grid, according to the operational requirements. The role of the charge/discharge controller is to charge and discharge the battery according to the renewable generation, electricity demand, and price of electricity. Meanwhile, the power grid is connected to the dairy farm and the battery. It supplies electricity when there is high demand and low renewable generation. The battery storage is used to satisfy the farm’s extra energy needs, a process commonly referred to as peak shaving. This involves using excess energy demands by utilizing stored power in the battery[41].
@@ 3.2 Data and Price Profile
For this study, two datasets were used: One dataset from Finland to train the algorithm and a second dataset collected from Ireland to evaluate the performance of the algorithm. The Finland dataset has information about the load demand from dairy farms, PV generation, wind generation, and electricity prices. The load data is collected from [42] and provides hourly electricity consumption over a year. Figure 2(a) demonstrates the monthly distribution of electricity demands for a dairy farm and PV generation and wind energy generated by the dairy farm throughout the year. The dataset consists of a dairy farm that has approximately 180 cows and has an estimated annual electricity usage of around 261 megawatt-hours (MWh). The PV and wind data was collected from the System Advisor Model (SAM) having a capacity of 20kW [43]. The Finland electricity price data was collected from a Helsinki electricity supply company [44]. This price data is dynamic and includes three different price levels [45]. The lowest rate is during off-peak hours, the standard rate applies for most of the day, and a higher peak rate is charged during the busiest hours. Specifically, the pricing is segmented into three time periods. The off-peak hours, with the lowest rate, are from 11 p.m. to 7 a.m. The standard rate applies during two intervals: from 8 a.m. to 4 p.m. and from 7 p.m. to 10 p.m. The peak rate, which is the highest, is charged between 5 p.m. and 7 p.m. Figure 2(b) shows how these electricity prices fluctuate over the day.
@@
The Ireland dataset includes data on farm load, PV generation, and electricity pricing. However, it lacks wind generation data as we were unable to find wind generation data for Ireland. The load consumption data, detailing electricity from the dairy farm over a year, was collected from a study on Irish dairy farms [46]. The PV generation data was collected from SAM[43] having a capacity of 20kW. The price data is collected from the Ireland electricity supply company Electric Ireland [47]. Figure 3 shows the Irish dairy farm energy consumption and photovoltaic (PV) energy generation and electricity price. This figure illustrates the variations in PV generation and electricity price, it also demonstrates the farm electricity demand patterns. The aim is to explore the relationship between energy consumption and PV generation, particularly in the Irish dairy farm context. Figure 3(a) specifically illustrates the monthly load demand and PV generation of the dairy farm over one year, while Figure 3(b) illustrates the price variations over the day.